# Navigating the Latent Space Dynamics of Neural Models

### v2

@@ Appendix D Additional implementation details
For the experiments in Figures 2,3, 13 we considered convolutional autoencoders with architecture specified in Table D. We train the models with Adam [25] with learning rate $45e-$ 4, linear step learning rate scheduler, and weight decay $41e-$ 4 for $500$ epochs. For the experiements in section 4.1, we use the pretrained autoencoder of [46]. We report in Table LABEL:tab:arch_autoencoderKL the architectural details and we refer to the paper for training details. For the experiments in section 4.2, we use the pretrained autoencoder of [15], considering the base model. We refer to the paper for training details. For computing attractors in experiments in Figures 2,3, 13, we compute the iterations $\mathbf{z}_{t+1}=f(\mathbf{z}_{t})$ until $\|f(\mathbf{z}_{t+1})-f(\mathbf{z}_{t})\|_{2}^{2}<1e-6$ or reaching $t=3000$. Similarly in experiments in Sections 4.1, 4.2, we compute the iterations $\mathbf{z}_{t+1}=f(\mathbf{z}_{t})$ until $\|f(\mathbf{z}_{t+1})-f(\mathbf{z}_{t})\|_{2}^{2}<1e-5$ or reaching $t=500$. All experiments are performed on a GPU NVIDIA 3080TI in Python code, using the PyTorch library [38].

### v3

@@ Appendix D Additional implementation details
For the experiments in Figures 2313 we considered convolutional autoencoders with architecture specified in Table D. We train the models with Adam [Kingma, 2014] with learning rate 5e45e-4, linear step learning rate scheduler, and weight decay 1e41e-4 for $500$ epochs. For the experiements in section 4.1, we use the pretrained autoencoder of [Rombach et al., 2022]. We report in Table LABEL:tab:arch_autoencoderKL the architectural details and we refer to the paper for training details. For the experiments in section 4.2, we use the pretrained autoencoder of [He et al., 2022], considering the base model. We refer to the paper for training details. For computing attractors in experiments in Figures 2,313, we compute the iterations $\mathbf{z}_{t+1}=f(\mathbf{z}_{t})$ until $\|f(\mathbf{z}_{t+1})-f(\mathbf{z}_{t})\|_{2}^{2}<1e-6$ or reaching $t=3000$. Similarly in experiments in Sections 4.14.2, we compute the iterations $\mathbf{z}_{t+1}=f(\mathbf{z}_{t})$ until $\|f(\mathbf{z}_{t+1})-f(\mathbf{z}_{t})\|_{2}^{2}<1e-5$ or reaching $t=500$. All experiments are performed on a GPU NVIDIA 3080TI in Python code, using the PyTorch library [Paszke et al., 2019].